# Efficient Frame Extraction: A Novel Approach Through Frame Similarity and Surgical Tool Tracking for Video Segmentation

### v1

@@ Abstract
The interest in leveraging Artificial Intelligence (AI) for surgical procedures to automate analysis has witnessed a significant surge in recent years. One of the primary tools for recording surgical procedures and conducting subsequent analyses, such as performance assessment, is through videos. However, these operative videos tend to be notably lengthy compared to other fields, spanning from thirty minutes to several hours, which poses a challenge for AI models to effectively learn from them. Despite this challenge, the foreseeable increase in the volume of such videos in the near future necessitates the development and implementation of innovative techniques to tackle this issue effectively. In this article, we propose a novel technique called Kinematics Adaptive Frame Recognition (KAFR) that can efficiently eliminate redundant frames to reduce dataset size and computation time while retaining useful frames to improve accuracy. Specifically, we compute the similarity between consecutive frames by tracking the movement of surgical tools. Our approach follows these steps: $i)$ Tracking phase: a YOLOv8 model is utilized to detect tools presented in the scene, $ii)$ Similarity phase: Similarities between consecutive frames are computed by estimating variation in the spatial positions and velocities of the tools, $iii$) Classification phase: A X3D CNN is trained to classify segmentation. We evaluate the effectiveness of our approach by analyzing datasets obtained through retrospective reviews of cases at two referral centers. The Gastrojejunostomy (GJ) dataset covers procedures performed between 2017 to 2021, while the Pancreaticojejunostomy (PJ) dataset spans from 2011 to 2022 at the same centers. In the GJ dataset, each robotic GJ video is segmented into six distinct phases. By adaptively selecting relevant frames, we achieve a tenfold reduction in the number of frames while improving accuracy by $4.32\%$ (from 0.749 to 0.7814). Our approach is also evaluated on the PJ dataset, demonstrating its efficacy with a fivefold reduction of data and a $2.05\%$ accuracy improvement (from 0.8801 to 0.8982). In addition, we also compare our approach with the state-of-the-art approaches to highlight its competitiveness in terms of performance and efficiency. Although we examined our approach on the GJ and PJ datasets for phase segmentation, this could also be applied to broader, more general datasets. Furthermore, KAFR can serve as a supplement to existing approaches, enhancing their performance by reducing redundant data while retaining key information, making it a valuable addition to other AI models.
@@ I Introduction
The growing application of AI in video analysis is primarily inspired by developments in modern Graphics Processing Units (GPUs) [1]. Traditional computer vision based video analysis for applications such as activity recognition, video summary, etc., faced challenges due to the higher dimensional nature of the data and the need to track both spatial and temporal features [2, 3]. However, the emergence of AI has enabled researchers to develop increasingly intricate and data-driven approach for interpreting video. This has led to considerable performance increases for a variety of applications, for example, object detection [4, 5, 6], action recognition [7, 8] and video segmentation [9, 10]. AI models can effectively learn complex patterns and temporal correlations directly from raw video data, revolutionizing the field and opening up new opportunities for applications in surveillance [11], autonomous vehicles [12] and healthcare [13].
@@
In the field of surgery, use of AI in analyzing surgical videos has a great potential [14, 15, 16], including automating activity recognition from video recordings [17, 18]. However, the duration of surgery videos might span from thirty minutes to several hours, posing a significant challenge for AI models, particularly in processing and analyzing lengthy sequential data effectively.
@@
Early approaches to handling this challenge leveraged statistical models designed for sequential data, such as conditional random field [19], and hidden Markov models [20]. However, due to the complex temporal relations among frames, these techniques exhibit limited representation capacities with predefined dependencies [21].
@@
Nonetheless, as deep learning technology evolved and gained prominence in the research community, a plethora of novel techniques have emerged, achieving impressive outcomes. When dealing with sequential data, models such as Gated Recurrent Unit GRU [22], integration of Long Shot-Term Memory (LSTM) and Convolutional Neural Netwoks (CNN) [23], and bi-directional LSTMs can be utilized [24].
@@
LSTM models, however, encounter the vanishing gradient problem when faced with long sequences of data. Alternatively, 3-D CNN models [3, 17] are favored for processing sequential data. This preference stems from their capacity to capture both spatial and temporal information simultaneously. A few notable examples illustrate this approach include a 3-D fully convolutional neural network (3DFCNN) that automatically encodes spatio-temporal patterns from these sequences, enabling action classification based on both spatial and temporal information. The method presents an end-to-end solution for human action recognition directly from raw depth image-sequences [25]. Additionally, CNN models have been effective in distinguishing between normal and Alzheimer’s disease patients utilizing an auto-encoder architecture and 3-D CNNs [26]. Furthermore, a fusion network that integrates the 3-D convolutional posture stream with the 2-D convolutional stream enhances the accuracy of identifying human actions [27]. However, compared to 2-D CNNs, 3-D CNNs can be more resource-intensive and challenging to train, especially when processing a large number of frames, due to their higher computational and memory needs.
@@
Vision transformers [2, 28, 29, 30] can handle longer sequences than CNN-LSTM models, but they require a large amount of data to train efficiently, which is often scarce and expensive in surgery.
@@
Since surgical videos are lengthy and complex, increasingly advanced techniques are needed to manage the intricate temporal dependencies. To address this issue, our approach involves fine-tuning by identifying key frames and removing non-essential ones, thereby reducing the total number of frames for training. Focusing on these key frames might help to train the network more efficiently by lowering computational costs, and speeding up the processing. Specifically, in this approach, i) background noise is filtered out while selectively preserving objects of interests, which are surgical tools in our work; ii) Adaptive 1 and Adaptive 2 methods are proposed to select key frames; iii) Distinct X3D CNN models are trained using the selected frames.
The primary contributions of this paper are:
@@
    Introduced the use of kinematics data (velocity and acceleration) with tool tracking for adaptively selecting key frames in video-based phase segmentation in surgery.
@@
    Developed and annotated a new and comprehensive Gastrojejunostomy (GJ) dataset.
@@
    Conducted extensive experiments on GJ and Pancreaticojejunostomy (PJ) datasets to validate the effectiveness of the proposed methods.
@@
    The source code is made available on GitHub (github.com) to facilitate reproduction of this work  [^1] https://github.com/leonlha/Kinematics-AFR.
The remainder of the article is organized as follows: In Section II, we discuss relevant approaches in applying Deep Learning in Surgery and selecting key frames. In addition, the proposed methods including Adaptive 1 and Adaptive 2 are introduced in Section III. The results of experiments and the extension of the experiments are presented in Sections V, and VI. Moreover, the comparison with state-of-the-art methods is discussed in Section VII, followed by discussions in Section VIII. Finally, we conclude our work in Section IX.
@@ II-A Surgery Datasets
One of the earliest surgical datasets, Cholec80, contains 80 videos of laparoscopic cholecystectomy surgeries performed by 13 surgeons. The dataset is divided into a training set (40 videos) and a testing set (40 videos), with the videos segmented into 7 phases. A subset of Cholec80, the Cholec51 dataset, includes 51 videos. However, our datasets differ as they focus on the reconstruction portion of the robotic pancreaticodudenectomy procedure, namely, the Gastrojejunostomy and Pancreaticojejunostomy anastomoses.
Additional datasets include the M2CAI16 Workflow Challenge, which consists of 41 laparoscopic cholecystectomy videos (27 for training and 14 for testing), segmented into 8 phases. The CATARACTS dataset, with 50 videos, features 19 phase categories. We collected a similar number of videos (GJ with 42 videos and PJ with 100 videos).
The JHU-ISI Gesture and Skill Assessment Working Set (JIGSAWS) contains 39 videos performed by 8 users, with each video recorded approximately five times. While JIGSAWS is frequently used in surgical skill assessment research, it focuses on simulation tasks, whereas our datasets were recorded in operating rooms during actual surgeries.
@@
Cholec80 [31] 80 7 Cholecystectomy procedures
@@
Cholec51 [31] 51 7 A subset of Cholec80
@@
 [32] 50 19 Cataract procedures from
@@
M2cai16 [33] 41 8 Minimally invasive surgical
@@
JIGSAWS [34] 17 4 Simulation procedures from
@@
PJ [3] 100 6 Pancreaticojejunostomy
GJ
42 6 Gastrojejunostomy (our)
The PJ dataset, which we collected previously, comprises 100 videos segmented into 6 phases of Pancreaticojejunostomy. We use this dataset to provide a comparative baseline for our current research.
@@ II-B Key Frame Recognition
As mentioned in the previous section, surgical videos tend to be longer than those in other fields, which poses a challenge for effectively using deep learning methods. A non-trivial approach entails significantly reducing the number of frames, for example, through Uniform Frame Sampling (UFS) which skips frames at regular intervals. Early attempts for key frame extraction relied on shot boundary-based algorithms [35]. Essentially, this technique utilizes the first or middle frame of each shot as the key frame after shot boundary detection. Although shot boundary-based approaches are simple to apply and generalize, the extracted key frame cannot fully capture the visual content. Another example used a Long Short-Term Memory (LSTM) network linked to the output of the underlying CNN [36]. However, this framework obtains sixteen sample frames evenly split with an eight-frame stride from the full-length videos as the video representation. Although straightforward, this technique disregards the significance of consecutive frames, implying that certain frames hold significantly more importance than others.
One solution involves computing the similarity between frames and discarding those that exceed a certain threshold. For instance, according to Savran et al. [37], frames are compared using measures like the Mean Squared Error (MSE) on human activity datasets, i.e., KTH and UCF-101. When high similarity between consecutive frames is observed, indicating minimal changes in the action, these frames are considered redundant and removed. By discarding frames that exceed a predefined similarity threshold, the technique effectively filters out frames that do not provide new or significant information to the video’s action. However, strategies that are effective in other fields may not be applicable in surgery. For example, tools for minimally invasive surgery, such as scissors and forceps (which need to be tracked), often occupy a small portion of the screen and are overshadowed by background noise. It is worth noting that human organs are constantly in motion, presenting significant challenges that render approaches like background subtraction, optical flow, and even depth estimation, as well as non-deep learning methods such as MSE and Structural Similarity Index (SSIM) (which use all pixels of the image), less efficient [37, 38, 39, 40].
@@ III-B1 Adaptive 1
In a video sequence of $n$ frames $\{x_{1},x_{2},...,x_{n}\}$, a threshold $d$ determines the distance measure $D(x_{i},x_{j})$ used to classify a pair of frames $(x_{i},x_{j})$ as key frames or similar frames. Key frame pairs are denoted as $P_{\text{key}}$. Frames $x_{i}$ and $x_{j}$ are considered key frames if their distance is less than or equal to $d$. The frames between the two key frames are similar and denoted by $P_{\text{similar}}$. The search for key frame pairs start from first frame until the last frame is encountered. Thus, we define a set of all key (K) pairs as follows:
$K(d)=\{(x_{i},x_{j})\mid D(x_{i},x_{j})\leq d,(x_{i},x_{j})\in P_{key}\},$ (1)
Assuming $S$ is a set of points (one pixel for each point) of a frame and $s$ is a subset of $S$, the distance $D(x_{i},x_{j})$ are defined as
@@
for constant $\beta_{d}>0$ and $\epsilon$ is a small number introduced to avoid division by zero.
Figure 2: Illustration of Adaptive Frame Recognition. The subset $s$ consists of two tools, $M$ and $N$ (represented as circle and diamond shapes), emerging from a noisy environment (which might include organs of the abdominal cavity, other surgical tools, and noise, represented by other shapes). The tools are tracked across frames. The frame index $k$ is incremented one at a time until a designated threshold $d$ is reached.
III-B2 Adaptive 2
@@
$K(d)=\{(x_{i},x_{j})\mid V(x_{i},x_{j})\leq d,(x_{i},x_{j})\in P_{key}\},$ (4)
Assuming $S$ is a set of points of a frame and $s$ is a subset $S$, the variation of velocity $V(x_{i},x_{j})$ is denoted as
$V(x_{i},x_{j})=f\left(\sum\limits_{\begin{subarray}{c}s\in S\end{subarray}}\sum_{k=i+1}^{j}\left|V_{\text{s}}(x_{i})-V_{\text{s}}(x_{k})\right|\right),$ (5)
@@
for constant $\beta_{v}>0$ and $\epsilon$ is a small number introduced to avoid division by zero.
III-C Object Detection
As mention earlier, the detection of surgical tools is crucial. Therefore, we propose using the YOLOv8 neural network, the latest advancement in the “You Only Look Once” (YOLO) [41]. It is well-known for its real-time object detection capabilities, outperforming those of fast and faster Region CNN (RCNN) [42] for tool detection (See Figure 3). While YOLO comes pre-trained on general datasets like COCO [43] or ImageNet [44], its direct application in surgical video analysis may be suboptimal. Surgical instruments and operating environments possess unique characteristics, necessitating fine-tuning of the model for optimal performance.
Figure 3: Tools auto-detection using Yolo.
We used a dataset comprising 28 surgical videos and employed the LabelMe [45] tool to generate annotations, yielding 917 annotated images (See Figure 4). To the best of our knowledge, there were not any other datasets containing the same set of surgical tools as those in the GJ and PJ datasets. Therefore, we used a portion of the GJ videos for tool segmentation. Although the same GJ dataset was used for both tasks–recognizing individual surgical tools and phase segmentation–these tasks are fundamentally different and should not interfere with each other’s performance.
In addition, LabelMe automatically labels any part of an image that is not explicitly masked by the user as “background”. This classification is crucial for training detection models because it trains them to distinguish between objects of interest and irrelevant areas.
@@ III-C Object Detection
YOLO framework splits the input image into a grid of $S\times S$ cells. Each cell generates candidate bounding boxes and confidence levels for instrument or tools presence. The confidence metric Conf(Tool) is determined by the binary indicator Pr(Tool) and is defined as:
@@
For more information on automated tool tracking, please refer to our other study [46].
III-D Surgical Phase Segmentation
III-D1 X3D Convolutional Neural Networks
Figure 5: Render X3D model architecture. (A) The X3D network extends a 2-D ResNet (res) one axis at a time across various axes, including temporal duration $X_{t}$, frame rate $X_{T}$, spatial resolution $X_{s}$, width $X_{w}$, bottleneck width $X_{b}$, and depth $X_{d}$. (B) Details of the 2-D ResNet demonstrate residual learning (Adapted from [47] and [48]).
Surgical phase segmentation is an important aspect of computer-assisted surgery and medical image analysis, involving the automated partitioning of surgical procedures into distinct phases or stages using video or image data. The primary goal of surgical phase segmentation is to comprehend the temporal progression of surgeries, facilitating various clinical applications such as workflow analysis and skill assessment [49, 50].
To identify these phases, we utilized the X3D architecture, a highly efficient 3-D CNN designed for activity recognition and video classification (refer to Figure 5). X3D is an extension of the 2-D ResNet model originally designed for image classification, using residual learning to preserve essential information as it propagates through layers [47, 48, 51].
The previous research [3] trained the X3D model based on a frame-by-frame approach, where, at any given time (e.g., at time $t_{i}$), the model was fed with $k$ frames, starting from frame $i$ and moving backward $k-1$ indices, as illustrated in the upper part of Figure 6. In their approach, frames were sampled at a fixed time interval. However, this research adopts a different strategy: we select only the key frames that capture significant tool movements. The lower part of the figure illustrates this difference. For example, as the tool moves with increasing velocity, the previous approach selects frames without accounting for this critical information. In contrast, our method selects only the key frames, such as the beginning and end frames, which not only dramatically reduce the size of the dataset but also signify moments that are likely to be crucial for classifying the phase.
Figure 6: Distinction between previous and proposed approaches. (A) Frame-by-Frame Level Phase Segmentation. (B) Uniform Frame Sampling vs. Kinematics Adaptive Frame Recognition.
III-D2 Two-stream Convolutional Networks
We integrated optical flow (OF) into our approach for video action recognition because of its critical role in capturing motion dynamics and temporal changes. Optical flow estimation involves computing angle and magnitude shifts for each pixel across consecutive video frames. The Farneback [52] algorithm is a common technique for dense optical flow computation, that utilizes polynomial expansion and iterative field estimation. By solving a quadratic polynomial equation between two consecutive frames $I(x,y,t)$ and $I(x+\Delta x,y+\Delta y,t+\Delta t)$, optical flow $(u,v)$ at each pixel is estimated to obtain displacement fields.

### v3

@@ Abstract
The interest in leveraging AI for surgical procedures to automate analysis has witnessed a significant surge in recent years. One of the primary tools for recording surgical procedures and conducting subsequent analyses, such as performance assessment, is through videos. However, these operating room (OR) videos tend to be notably lengthy compared to other fields, spanning from thirty minutes to more than one hour, which poses a challenge for AI models to effectively learn from them. Despite this challenge, the foreseeable increase in the volume of such videos in the near future necessitates the development and implementation of innovative techniques to tackle this issue effectively. In this article, we propose a novel technique called Kinematics Adaptive Frame Recognition (KAFR) that can efficiently eliminate redundant frames to reduce dataset size and computation time while retaining useful frames to improve accuracy. Specifically, we compute the similarity between consecutive frames by tracking the movement of surgical tools. Our approach follows these steps: $i)$ Tracking phase: a YOLOv8 model is utilized to detect tools presented in the scene, $ii)$ Similarity phase: Similarities between consecutive frames are computed by estimating variation in the spatial positions and velocities of the tools, $iii$) Classification phase: A X3D CNN is trained to classify segmentation. We evaluate the effectiveness of our approach by analyzing [..111removed: a dataset obtained through a retrospective review of cases spanning from ] datasets obtained through retrospective reviews of cases at two referral centers. The Gastrojejunostomy (GJ) dataset covers procedures performed between 2017 to 2021 [..222removed: at two referral institutes. Within this ] , while the Pancreaticoduodenectomy (PJ) dataset spans from 2011 to 2022 at the same centers. In the GJ dataset, each robotic [..333removed: Gastrojejunostomy (GJ ) video was ] GJ video is segmented into six distinct phases. By adaptively selecting relevant frames, we achieve a tenfold reduction in the number of frames while increasing accuracy by three percent. Our approach is also evaluated on the PJ dataset, demonstrating its efficacy. In addition, we also compare our approach with [..444removed: a ] the state-of-the-art [..555removed: approach to demonstrate its superiority ] approaches to highlight its competitive in terms of performance and efficiency. Although we examined our approach on the GJ [..666removed: dataset ] and PJ datasets for phase segmentation, this could also be applied to broader, more general datasets. Furthermore, KAFR can be used as a supplement to existing approaches, enhancing their performance by reducing redundant data while retaining key information, making it a valuable addition to other AI models. Our source code is ready for replication at https://github.com/leonlha/Kinematics-AFR.
This paper includes surgical images that may be distressing to some readers.
@@ I Introduction
The growing application of AI in video analysis is primarily inspired by developments in modern Graphics Processing Units (GPUs) [phong2022pso]. Traditionally, video analysis presented challenges due to their high dimension, temporal dynamics and visual variation. However, the emergence of AI has enabled researchers to develop increasingly intricate and data-driven approach for interpreting video. This has led to considerable performance increases for a variety of applications, for example, object detection [liu2023yolo, liang2015recurrent, felzenszwalb2009object], action recognition [sun2024k, phong2023pattern] and video segmentation [garg2024human, grammatikopoulou2024spatio]. AI models can effectively learn complex patterns and temporal correlations directly from raw video data, revolutionizing the field and opening up new opportunities for applications in surveillance [li2020abnormal], autonomous vehicles [gao2024human] and healthcare [IJMESD400].
@@
In the field of surgery, there is an increasing trend towards leveraging advancements in AI, particularly for automating the analysis of surgical processes extracted from video recordings  [demir2023deep]. However, the duration of surgery videos might spans from thirty minutes to several hours, posing a significant challenge for AI models [..777removed: . Initially, statistical models ] , particularly in processing and analyzing lengthy sequential data effectively.
@@
Early approaches to handling this challenge leveraged statistical models designed for sequential data, such as conditional random field [quellec2014real] , and hidden markov models [padoy2008line] [..888removed: were utilized ]. However, due to the complex temporal relations among frames, these techniques exhibit limited representation capacities with predefined dependencies [jin2017sv].
@@
Nonetheless, as deep learning technology evolved and gained prominence in the research community, a plethora of novel techniques have emerged, achieving impressive outcomes. When dealing with sequential data, models such as Gated Recurrent Unit GRU [valipour2017recurrent], integration of Long Shot-Term Memory (LSTM) and Convolutional Neural Netwoks (CNN) [..999removed: can be utilized [jalal2021deep]. Beyond-Short-Snippets [yue2015beyond] aimed to extract feature maps from 2-D ConvNets and integrate them with LSTMs for video predictions. The Gated Recurrent Unit (GRU) was introduced with the aim of efficiently handling temporal data in the context of online segmentation [valipour2017recurrent]. Alternative approaches have also utilized ]  [jalal2021deep], and bi-directional LSTMs [..101010removed: , which consist of two distinct LSTMs, to explore both forward and backward temporal information ] can be utilized [he2021db].
@@
[..111111removed: Though, LSTM models ] LSTM models, however, encounter the vanishing gradient problem when faced with long sequences of data. Alternatively, 3-D CNN models [al2024development, hegde2024automated] are favored for processing sequential data. This preference stems from their capacity to capture both spatial and temporal information simultaneously. A few notable examples illustrate this approach: Fractioned Adjacent Spatial and Temporal (FAST) 3-D convolutions are used to extract spatio-temporal patterns from 3-D videos, expanding on 2-D image convolutions [stergiou2019spatio]. CNN models have effectively distinguished between normal and Alzheimer’s disease patients utilizing an auto-encoder architecture and 3-D CNNs [hosseini2016alzheimer]. A fusion network that integrates the 3-D convolutional posture stream with the 2-D convolutional stream enhances the accuracy of human action recognition [huang2018human].
However, 3-D CNN can become bulky and challenging to train when using a large number of frames.
@@
Vision transformers [nguyen2023video, han2023flatten, liu2021swin, arnab2021vivit] can handle longer sequences than CNN-LSTM models, but they require a large amount of data to train efficiently, which is often scarce and expensive in surgery.
@@
[..121212removed: Hence, a ] Since surgical videos are lengthy and complex, increasingly advanced techniques are needed to manage the intricate temporal dependencies. To address this issue, our approach involves fine-tuning by identifying key frames and removing non-essential frames, thereby reducing the number of frames for training. Focusing on these key frames might help to train the network more efficiently by lowering computational costs, and speeding up processing. Specifically, in this approach, (1) background noise is filtered out while selectively preserving crucial surgical tools. (2) Adaptive 1 and Adaptive 2 methods are proposed to select most important frames. (3) Distinct X3D CNN models are used to train on the selected frames and corresponding optical flow data before ensembling.
The primary contributions of this paper include:
@@
    Introducing the use of kinematics data (velocity and acceleration) for adaptively selecting key frames in video-based phase segmentation in surgery, an crucial approach in this field.
@@
    Presenting a novel adaptive method that achieves state-of-the-art results.
@@
    Developing and annotating a new and comprehensive GJ dataset
@@
    Conducting extensive experiments on GJ dataset and PJ dataset to validate the effectiveness of the proposed methods.
  5. 5.
    The source code for our proposed methods is available for reproduction on GitHub at github.com
The rest of the article is organized as follows: In Section II, we discuss relevant approaches in applying Deep Learning in Surgery and selecting key frames. In addition, the proposed methods including Adaptive 1 and Adaptive 2 are introduced in Section III. The results of experiments and the extension of the experiments are presented in Sections V, and VI. Finally, we conclude our work in Section VIII.
@@ II-A Surgery Datasets
Cholec80 [twinanda2016endonet] 80 7 Cholecystectomy procedures
@@
Cholec51 [twinanda2016endonet] 51 7 A subset of Cholec80
@@
 [ALHAJJ201924] 50 19 Cataract procedures from
@@
M2cai16 [cadene2016m2cai] 41 8 Minimally invasive surgical
@@
JIGSAWS [gao2014jhu] 17 4 Simulation procedures from
@@
PJ [al2024development] 100 6 Pancreaticojejunostomy
GJ 42 6 Gastrojejunostomy
(this work)
One of the earliest surgical datasets, Cholec80, contains 80 videos of cholecystectomy surgeries performed by 13 surgeons. The dataset is divided into a training set (40 videos) and a testing set (40 videos), with the videos segmented into 7 phases. A subset of Cholec80, the Cholec51 dataset, includes 51 videos. However, our datasets differ as they focus on the procedures of Gastrojejunostomy and Pancreaticoduodenectomy, rather than cholecystectomy.
Additional datasets include the M2CAI16 Workflow Challenge dataset, which consists of 41 laparoscopic cholecystectomy videos (27 for training and 14 for testing), segmented into 8 phases. The CATARACTS dataset, with 50 videos, features 19 phase categories. We collected a similar number of videos (GJ with 42 videos and PJ with 100 videos).
The JHU-ISI Gesture and Skill Assessment Working Set (JIGSAWS) contains 39 videos performed by 8 users, with each video recorded approximately five times. While JIGSAWS is frequently used in surgical skill assessment research, it focuses on simulation tasks, whereas our dataset was recorded in an operating room during actual surgeries.
The PJ dataset, which we collected previously, comprises 100 videos segmented into 6 phases of Pancreaticoduodenectomy procedures. We use this dataset to provide a comparative baseline for our current research.
@@ II-B Key Frame Recognition
As mentioned in the previous section, surgical videos tend to be longer than those in other fields, which poses a challenge for effectively using deep learning methods. A non-trivial approach entails significantly reducing the number of frames, for example, through Uniform Frame Sampling (UFS) which skips frames at regular intervals. Early attempts for key frame extraction relied on shot boundary-based algorithms [boreczky1996comparison]. Essentially, this technique utilizes the first or middle frame of each shot as the key frame after shot boundary detection. Although shot boundary-based approaches are simple to apply and generalize, the extracted key frame cannot fully capture the visual content. Another example used a Long Short-Term Memory (LSTM) network linked to the output of the underlying CNN [donahue2015long]. However, this framework obtains sixteen sample frames evenly split with an eight-frame stride from the full-length videos as the video representation. Although straightforward, this technique disregards the significance of consecutive frames, implying that certain frames hold significantly more importance than others.
One solution involves computing the similarity between frames and discarding those that exceed a certain threshold. However, strategies that are effective in other fields may not be applicable in surgery. For example, tools for minimally invasive surgery, such as scissors and forceps, often occupy a small portion of the screen and are overshadowed by background noise. It is worth noting that human organs are constantly in motion, presenting significant challenges that render approaches like background subtraction, optical flow, and even depth estimation, as well as non-deep learning methods such as Mean Square Error (MSE) and SSIM (which use all pixels of the image), less efficient [savran2023novel, yang2021fast, zhong2020key, cucchiara2003detecting].
[..131313removed: Our approach is based on a fine-tune process of filtering out background noise while selectively preserving those that play a crucial role in identifying key frames. By prioritizing the retention of only the most relevant elements, we aim to improve the efficiency of our method. This strategic refinement not only reduces the computational demands placed on our system but also enhances the precision and accuracy of our results. In essence, our strategy optimizes the balance between computational efficiency and analytical reliability, hence increasing the overall technical efficacy of our approach. ]
[..141414removed: The primary contributions of this paper include: ] [..151515removed: Introducing the use of kinematics data (velocity and acceleration)for adaptively selecting key frames in video-based phasesegmentation in surgery, an crucial approach in this field. ] [..161616removed: Presenting a novel adaptive method that achieves state-of-the-art results. ] [..171717removed: Developing and annotating a new and comprehensive GJ dataset ] [..181818removed: Conducting extensive experiments on GJ dataset to validate the effectiveness of the proposed methods. ] [..191919removed: The source code for our proposed methods is available for reproduction on GitHub at github. com ]
@@ III-B1 Adaptive 1
In a video sequence of frames $\{x_{1},x_{2},...,x_{n}\}$, a threshold $d$ determines the distance measure $D(x_{i},x_{j})$ used to classify a pair of frame $(x_{i},x_{j})$ as key frame or similar frame. Key frame pairs are denoted as $P_{\text{key}}$. Frames $x_{i}$ and $x_{j}$ are considered key frames if their distance is less than or equal to $d$. The frames between the two key frames are similar $P_{\text{similar}}$. The search for key frame pairs start from first frame until the last frame is encountered. Thus, we define a set of all key (K) pairs as follows:
$K(d)=\{(i,j)\mid D(x_{i},x_{j})\leq d,(i,j)\in P_{key}\}{\color[rgb]{0,0,1}\sf,}$ (1)
Assuming $S$ is a set of points of a frame and $s$ is a subset of $S$, the distance $D(x_{i},x_{j})$ are defined as
@@
for constant $\beta_{d}>0$ and $\epsilon$ is introduced to avoid division by zero.
Figure 2: Illustration of Adaptive Frame Recognition. The subset $s$ consists of two tools, $M$ and $N$ (represented as circle and diamond shapes), [..232323removed: emerged ] emerging from a noisy environment (which might include organs of the abdominal cavity, other surgical tools, and noise, represented by other shapes). The tools are tracked across frames. The frame index $k$ is incremented one at a time until a designated threshold $d$ is [..242424removed: achieved ] reached.
III-C [..252525removed: Alternative KAFR ]
III-B1 Adaptive 2
@@
$K(d)=\{(i,j)\mid V(x_{i},x_{j})\leq d,(i,j)\in P_{key}\}{\color[rgb]{0,0,1}\sf,}$ (4)
Assuming $S$ is a set of points of a frame and $s$ is a subset $S$. The variation of velocity $V(x_{i},x_{j})$ is denoted as
$V(x_{i},x_{j})=f\left(\sum\limits_{\begin{subarray}{c}s\in S\end{subarray}}\sum_{k=i+1}^{j}\left|V_{\text{s}}(x_{i})-V_{\text{s}}(x_{k})\right|\right){\color[rgb]{0,0,1}\sf,}$ (5)
@@
for constant $\beta_{v}>0$ and $\epsilon$ is introduced to avoid division by zero.
IV [..272727removed: Experiments ]
III-A [..292929removed: Benchmark Dataset ] Object Detection
[..313131removed: Pancreatic cancer is projected to become the second leading cause of cancer-related deaths in the USA within the next two to three decades [mizrahi2020pancreatic]. Robotic Pancreaticoduodenectomy (also known as the Whipple procedure) surgery, which is a key procedure for treating pancreatic cancer, involves two main phases: Resection and Reconstruction. The Reconstruction phase includes three anastomoses: Hepaticojejunostomy, Pancreaticojejunostomy, and Gastrojejunostomy. ]
[..323232removed: A retrospective case review spanning from 2017 to 2021 was conducted at two prominent referral centers, UT Southwestern and University of Pittsburgh Medical Center. During this review, each robotic GJ video was segmented into six specific tasks: 1.1 Stay suture, 1.2 Inner running suture, 1.3 Enterotomy, 2.2 Inner running suture, 3.1 Inner Layer of Connell, and 4.1 Outer Layer of Connell, along with idle time. These videos were annotated by experts using a custom annotation framework. ]
[..333333removed: Of the 42 videos included, 30 were used for training and 12 for validation, with all frames extracted (six frames/second) and annotated, resulting in a dataset of 46,903 images for training and 22,647 images for validation. ]
III-B [..343434removed: Distribution of Tools ]
[..363636removed: Figure 7 displays the distribution of the number of frames for each class ID corresponding to the sixteen objects listed in Table II. We used YOLOV8 [redmon2016you], as described in Section III-B, to extract and track surgical instruments. Each instrument is segmented into three parts, each assigned a unique ID: the shaft, wrist, and jaw, except for the “laptool”. It’s important to note that in our dataset, instrument number two (“instrument2”) is identical to instrument number three (“instrument3”) because they are both needle drivers; therefore, they are combined under the same object category. ] [..373737removed: hbtp! ] [..383838removed: Distribution of Class IDs. Total of sixteen objects are tracked. ]
III-B [..393939removed: Object Detection ]
[..414141removed: The research utilized the ] As mention earlier, the detection of surgical tools is crucial. Therefore, we propose using the YOLOv8 neural network, the latest advancement in the “You Only Look Once” (YOLO) versions and is well-known for its real-time object detection capabilities, which outperform those of fast and faster Region CNN (RCNN) [girshick2014rich] for tool detection (See Figure 3). While YOLO comes pre-trained on general datasets like COCO [lin2014microsoft] or ImageNet [deng2009imagenet], its direct application in surgical video analysis may be suboptimal. Surgical instruments and operating environments possess unique characteristics, necessitating fine-tuning of the model for optimal performance.
Figure 3: [..444444removed: LabelMe for surgical tools annotation ] Tools auto-detection using Yolo.
We used a dataset comprises of 28 surgical videos and employed the LabelMe [russell2008labelme] tool to generate annotations, yielding 917 annotated images [..454545removed: . ] (See Figure 4). To the best of our knowledge, there were no other datasets containing the same set of surgical tools as those in the GJ and PJ datasets. Therefore, we used a portion of the GJ videos for tool segmentation. Although the recognition of individual surgical tools is distinct from phase segmentation, these two tasks are fundamentally different and should not affect each other’s performance. In addition, LabelMe automatically labels any part of an image that is not explicitly masked by the user as “background”. This classification is crucial for training detection models because it trains them to distinguish between objects of interest and irrelevant areas.
@@
YOLO framework splits the input image into a grid of $S\times S$ cells. Each cell generates candidate bounding boxes and confidence levels for instrument or tools presence. The confidence metric Conf(Tool) is determined by the binary indicator Pr(Tool) and is defined as: [..464646removed: tbp! ] [..474747removed: Tools auto-detection using Yolo. ]
@@
III-B Surgical Phase Segmentation
III-B1 X3D Convolutional Neural Networks
Figure 5: Render X3D model architecture. (A) The X3D network extends a 2-D ResNet (res) one axis at a time across various axes, including temporal duration $X_{t}$, frame rate $X_{T}$, spatial resolution $X_{s}$, width $X_{w}$, bottleneck width $X_{b}$, and depth $X_{d}$. (B) Details of the 2-D ResNet demonstrate residual learning (Adapted from [feichtenhofer2020x3d] and [he2016deep]).
Surgical phase segmentation is an important aspect of computer-assisted surgery and medical image analysis, involving the automated partitioning of surgical procedures into distinct phases or stages using video or image data. The primary goal of surgical phase segmentation is to comprehend the temporal progression of surgeries, facilitating various clinical applications such as workflow analysis and skill assessment [tran2017phase].
To identify these phases, we utilized the X3D architecture, a highly efficient 3-D CNN designed for activity recognition and video classification (refer to Figure 5). X3D is an extension of the 2-D ResNet model originally designed for image classification, using residual learning to preserve essential information as it passes through layers.
III-B2 Two-stream Convolutional Networks
We integrated optical flow (OF) into our approach for video action recognition because of its critical role in capturing motion dynamics and temporal changes. Optical flow estimation involves computing angle and magnitude shifts for each pixel across consecutive video frames. The Farneback [farneback2003two] algorithm is a common technique for dense optical flow computation, that utilizes polynomial expansion and iterative field estimation. By solving a quadratic polynomial equation between two consecutive frames $I(x,y,t)$ and $I(x+\Delta x,y+\Delta y,t+\Delta t)$, optical flow $(u,v)$ at each pixel is estimated to obtain displacement fields.